OpenStudy (anonymous):

x varies inversely as y^2, and x=4 when y=10. find x when y=2

OpenStudy (anonymous):

Since Inversely proportionals is given: \[x = \frac{k}{y^2}\] Where k is any constant.. Now you have given x = 4 and y = 10 Plug in the values and find k first..

OpenStudy (anonymous):

so like this 4 = k/2^2

OpenStudy (anonymous):

Yes but at first y is given 10.. So use y = 10..

OpenStudy (anonymous):

so 400

OpenStudy (anonymous):

Yes you are right.. Now you got k = 400 (remember this value).. Now, we have to find x but y is given as 2.. Can you find x ?? k = 400 and y = 2

OpenStudy (anonymous):

100

OpenStudy (anonymous):

Yes absolutely right... Good Job..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

i just did help setting it up

OpenStudy (anonymous):

I have one more like that I am gonna try it can you tell me if I am right

OpenStudy (anonymous):

See the difference here if the question says x varies directly as y squared then the equation would be: \[x = ky^2\] In short: Inversely means in denominator. Directly means in numerator.. And note that the constant k will always be in Numerator..

OpenStudy (anonymous):

okay thanks again

OpenStudy (anonymous):

Take your time try it once and I will check if it is right or not..

OpenStudy (anonymous):

x varies inversely as v, and x = 48 when v = 8. Find x when v = 64 so I got x = 6

OpenStudy (anonymous):

What you got for k ??

OpenStudy (anonymous):

384

OpenStudy (anonymous):

Yes dear..

OpenStudy (anonymous):

384/64 = 6

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Now do you understand how to solve them??

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

So, Best Of Luck.. Generally I do not give medals easily but this one for you for doing the Second Problem firstly on your own and secondly absolutely right... Ha ha ha ha..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

it has just been a long time for me I have been out of school for over 25 years and I am trying to get my high school diploma by myself online

OpenStudy (anonymous):

Your age???

OpenStudy (anonymous):

41

OpenStudy (anonymous):

i quit school when I was 16

OpenStudy (anonymous):

never had freshman credits

OpenStudy (anonymous):

O My God, can you imagine I am exactly 20 years younger than you.. I am just 21 now.. Doesn't matter.. Forget the past, Live in Present and Think about making the Future wonderful.. Best of Luck dear.. I pray, you will have your diploma at the earliest...

OpenStudy (anonymous):

thank you so much I raised my kids they are all in college now it is time for myself

OpenStudy (anonymous):

Yes Best of Luck for today tomorrow and for future...